OpenStudy (anonymous):

A recipe calls for 10 cups of oatmeal. Benjamin wants to make fewer servings than the recipe makes. Which expression represents how many cups of oatmeal Benjamin should use? A. 0.8 10 B. 1.8 10 C. 0.2 10 D. 1.2 10

OpenStudy (mathstudent55):

How many fewer servings does he want to make? The number is missing in your problem statement above.

OpenStudy (anonymous):

how any fewer servings does he want to make, that's the only number your missing

OpenStudy (anonymous):

do you know what 1/5 is in decimal form?

OpenStudy (anonymous):

@homeworkblah

OpenStudy (anonymous):

1/5 = 2/10 what is 2/10 in decimal form @homeworkblah?

OpenStudy (anonymous):

2/10 = 0.2 @homeworkblah

OpenStudy (anonymous):

to solve this we simply do it like this, 1/5 x 10 remember that 1/5 = 0.2 0.2 x 10 = ? Can you figure that out @homeworkblah?

OpenStudy (anonymous):

He wants to make 4/5 the servings. 4/5 as a decimal is 0.8 So A, 0.8 * 10 is the answer